OpenStudy (anonymous):

find the inverse laplace of 1/(s(s-3))

OpenStudy (experimentx):

\[ {1 \over s (s-3)} = {1 \over -3}\left ( {1 \over s} - {1 \over s- 3} \right)\] take inverse Laplace transform.

OpenStudy (anonymous):

what did you do?

OpenStudy (abb0t):

\[\frac{ 1 }{ 3 }e ^{3t}\]

OpenStudy (anonymous):

okay so you split it up right?

OpenStudy (experimentx):

yep!! http://en.wikipedia.org/wiki/Heaviside_cover-up_method

OpenStudy (anonymous):

i dont understand so 1/s + 1/(s-3)

OpenStudy (abb0t):

\[\frac{ 1 }{ 3 }(1-e ^{3t})\]

OpenStudy (anonymous):

i dont need the answer haha

OpenStudy (anonymous):

And giving answer is against the Code of Conduct of this site!

OpenStudy (abb0t):

You get it in a form so that you can use the laplace formulas.

OpenStudy (anonymous):

okay so can i use partial fractions

OpenStudy (anonymous):

like 1/3(s-3) - 1/3s

OpenStudy (unklerhaukus):

Partial fractions \[\frac1{s(s-3)}=\frac{A}{s}+\frac{B}{s-3}\] \[1=A(s-3)+B(s)\] when \(s=0\) \[1=-3A\qquad \implies\qquad A=-\frac13\] when \(s=3\) \[1=3B\qquad \implies\qquad B=\frac13\] \[\frac1{s(s-3)}=\frac{-1/3}{s}+\frac{1/3}{s-3}=\frac13\left(\frac1{s-3}-\frac{1}{s}\right)\]

OpenStudy (anonymous):

okay where do i go from there?

OpenStudy (anonymous):

1/s is just t

OpenStudy (anonymous):

i meant 1

OpenStudy (unklerhaukus):

now take the inverse laplace transform

OpenStudy (anonymous):

how do i take it for 1/3(s-3)

OpenStudy (anonymous):

oh its just e^455465

OpenStudy (unklerhaukus):

\[\mathcal L\Big\{\frac13\left(\frac1{s-3}-\frac{1}{s}\right)\Big\}\] \[=\frac13\left(\mathcal L\Big\{\frac1{s-3}\Big\}-\mathcal L\Big\{\frac{1}{s}\Big\}\right)\]

OpenStudy (anonymous):

1/3 * e^3t - 1/3

OpenStudy (unklerhaukus):

\[\huge\color{red}\checkmark\]